### v1

# FedCCRL: Federated Domain Generalization with Cross-Client Representation Learning

Abstract
Domain Generalization (DG) aims to train models that can effectively generalize to unseen domains. However, in the context of Federated Learning (FL), where clients collaboratively train a model without directly sharing their data, most existing DG algorithms are not directly applicable to the FL setting due to privacy constraints, as well as the limited data quantity and domain diversity at each client. To tackle these challenges, we propose FedCCRL, a novel federated domain generalization method that significantly improves the model’s ability to generalize to unseen domains without compromising privacy or incurring excessive computational and communication costs. Specifically, we adapt MixStyle to the federated setting to transfer domain-specific features, while AugMix is employed to perturb domain-invariant features. Furthermore, we leverage supervised contrastive loss for representation alignment and utilize Jensen-Shannon divergence to ensure consistent predictions between original and augmented samples. Extensive experimental results demonstrate that FedCCRL achieves the state-of-the-art performance on the PACS, OfficeHome, and miniDomainNet datasets across varying numbers of clients. Code is available at https://github.com/SanphouWang/FedCCRL.
1 Introduction
Domain Generalization (DG) [44] seeks to train models that can generalize effectively to unseen domains, a crucial task in real-world applications where the test data may come from domains that differ from the training data [26]. To achieve this, DG methods focus on two types of features: domain-invariant features and domain-specific features. Domain-invariant features (e.g., objects’ shape in image classification) are those that remain consistent across different domains and are essential for enabling a model to generalize to unseen domains. On the other hand, domain-specific features (e.g., color distribution of images in Figure 1) are unique to each domain and hinder generalization if overemphasized by the model.
Figure 1: Illustration of the typical setting in FL. Each client contains data from a unique domain, and the test domain (Photo) differs from all domains present on the clients.
To enhance models’ generalization ability, representation learning plays a vital role in existing DG methods [1, 25, 23, 17]. A key component of this is representation alignment, which focuses on aligning feature representations of samples across domains to ensure that domain-invariant features are effectively learned. This alignment helps models capture patterns that are shared across domains while minimizing reliance on domain-specific features. In addition, feature augmentation [43, 37, 11, 41] is another powerful strategy for DG. Methods like MixStyle [43] generate new samples by mixing instance-level statistics from different domains. This approach exposes the model to a broader range of examples, thereby enhancing its ability to capture domain-invariant features while mitigating reliance on domain-specific characteristics [22], such as lighting conditions or writing styles.
We refer to these methods as centralized DG methods, as they typically assume that all data is centrally stored and accessible. However, in real-world applications, data is often distributed across multiple clients, making direct data sharing impractical due to privacy concerns. This challenge is addressed by Federated Learning (FL) [19], where clients collaboratively train models without sharing their local datasets. In order to enhance the domain generalization ability of models in FL systems, Federated Domain Generalization (FDG) [18] has been receiving increasing attention. Figure 1 provides a typical FDG setting, where each client has access to data from a single domain and the target domain differs from those on all clients. However, in such FDG setting, the limited quantity of samples and the lack of domain diversity on each client complicates the direct application of centralized DG methods [2].
Most existing FDG methods [24, 39, 7] attempt to boost the model’s generalization capacity by aligning representation distributions and extracting domain-invariant features. However, due to the limited data quantity and domain diversity on individual clients, the performance of these methods often falls short. To mitigate these shortcomings, CCST [3] leverages AdaIN [10] for cross-client style transfer to increase domain diversity on each client. However, CCST introduces additional networks and requires the transmission of a large number of high-dimensional feature representations, resulting in significant computational overhead and transmission costs. Moreover, transmitting sample features poses a considerable privacy risk, as attackers could reconstruct highly realistic samples with these feature representations and the same decoder [21, 15].
To address the challenges in FDG, we propose FedCCRL—a novel method that significantly improves model generalization in the FL setting while maintaining privacy and incurring only negligible additional computational and communication costs. Specifically, FedCCRL combines cross-client feature augmentation, representation alignment, and prediction alignment. In the feature augmentation module, we adapt MixStyle [43] to the Federated Learning (FL) setting for cross-client transfer of domain-specific features, and we apply AugMix [8] to perturb domain-invariant features. For representation alignment, we utilize supervised contrastive loss [12] to align the representations of original and augmented samples. Lastly, prediction alignment leverages Jensen-Shannon divergence [20] to ensure consistency in predictions between original and augmented samples.
The key contributions of this work are outlined as follows:
  • We propose FedCCRL, a novel federated domain generalization framework that integrates cross-client feature augmentation, representation alignment, and prediction alignment to improve the model’s generalization capability in federated settings.
  • We adapt MixStyle to the federated learning setting, enabling cross-client style transfer while maintaining data privacy and only introducing minimal additional computational and communication costs
  • Extensive experiments on PACS, OfficeHome, and miniDomainNet datasets demonstrate that the proposed FedCCRL method achieves state-of-the-art (SOTA) FDG performance across FL systems of varying numbers of clients.
2 Related Work
This section reviews key research areas related to DG [44], with a focus on representation alignment, feature augmentation, and Federated Domain Generalization (FDG). These aspects are critical in addressing the challenges of domain shifts and ensuring robust model performance in federated learning.
2.1 Representation Alignment
One of the key techniques in representation learning for DG is representation alignment, where feature representations from different domains are aligned to minimize domain-specific variations. For instance, DANN [5, 4, 6] employs adversarial learning to align the representation distributions across domains by using a domain classifier to ensure that the learned features are domain-invariant. Another prominent approach is CORAL [28], which aligns the second-order statistics of source and target representation distributions to reduce domain shift. Additionally, MMD-based methods [29, 33, 34] utilize kernel-based distances to align representations across domains. These methods focus on reducing domain-specific discrepancies while maintaining discriminative power, helping models generalize better to unseen domains.
2.2 Feature Augmentation
Feature augmentation [31, 32, 36] has also emerged as an effective strategy for DG by generating diverse variations of the data to improve models’ robustness. One notable method is MixStyle [43], which generates new feature samples by mixing instance-level statistics (mean and variance) from different domains, effectively expanding the feature space and enabling models to generalize better across unseen domains. In addition to MixStyle, Mixup [38, 35] creates new data by linearly interpolating between two instances and their labels, using a weight sampled from a Beta distribution. Both methods are conceptually simple and computationally efficient, yet they have shown strong performance on popular benchmarks.
2.3 Federated Domain Generalization
Most existing FDG algorithms address these challenges by introducing regularization terms to enable the model to learn domain-invariant features or to ensure that feature representations across clients are aligned to a common distribution. For example, FedADG [39] utilizes federated adversarial learning to align distributions across clients to a reference distribution. FedSR [24] adopts an L2-norm regularizer and a conditional mutual information regularizer to align feature distributions across domains. In addition, FedIIR [7] takes a different approach by implicitly learning invariant relationships through gradient alignment, helping the model generalize better to OOD data. Moreover, GA [40] introduces novel global objective and optimizes it by dynamically calibrating the aggregation weights.
Although the aforementioned FDG algorithms have improved the model’s generalization ability to some extent, they do not directly address the issue posed by limited data volume and domain diversity on each client and fail to achieve the desired performance. To this end, CCST [3] employs a cross-client style transfer approach based on AdaIN [10] to generate samples with styles from other domains on the client side. However, CCST relies on a pre-trained VGG network [27] to extract feature representations and perform sample reconstruction, while also needing to transmit these high-dimensional representations, which results in significant additional communication and computational overhead. Furthermore, transmitting these feature representations violates privacy protection principles in FL, as attackers can easily intercept these representations and reconstruct the original samples with the same decoder [21, 15]. Additionally, the use of a pre-trained VGG network also partially violates the principles of domain generalization, as the target domain might be included in the pre-training dataset.
3 Method
In this section, we begin by introducing a standard FDG setting. We then describe the feature augmentation techniques utilized in FedCCRL, followed by a detailed explanation of the feature alignment regularization and the Jensen-Shannon divergence regularization. An overview of the proposed FedCCRL framework is provided in Figure 2, with the algorithmic details outlined in Algorithm 1.
Figure 2: Overview of FedCCRL. The client transmits local model parameters and sample statistics to the server, where these elements are aggregated and redistributed to each client. During the local training phase, feature augmentation, representation alignment and prediction alignment are applied to ensure the model learns to focus on domain-invariant features.
3.1 Preliminary
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input space and target space, respectively. In FDG, given $M$ source domains $S_{source}=\{S_{i}|i=1,2,\dots,M\}$, the data points in $S_{i}$ are independently and identically drawn from distribution $P_{i}(x,y)$, where $x\in\mathcal{X}$, $y\in\mathcal{Y}$. The joint distribution between each pair of domains are different, i.e., $P_{i}(x,y)\neq P_{j}(x,y)$ for $1\leq i\neq j\leq M$.
In this study, we adopt a commonly used data partitioning scheme in FDG, where the dataset on each client, denoted as $D_{i}$, originates exclusively from a single domain. Specifically, each source domain is partitioned into $K$ clients: $\cup_{j=k\cdot(i-1)+1}^{i\cdot k}D_{j}=S_{i}$, and $D_{q}\cap D_{p}=\emptyset$ for $1\leq q\neq p\leq K\cdot M$.
The goal of FDG is to enable these $K\cdot M$ clients to collaboratively train a model $f:\mathcal{X}\to\mathcal{Y}$, such that the model is capable of generalizing well and minimize the prediction error on a target domain $S_{\text{target}}$:
$\min_{f}\mathbb{E}_{(x,y)\sim S_{\text{target}}}\left[l(f(x),y)\right],$ (1)
where $l(\cdot)$ represents the loss function. Samples in $S_{\text{target}}$ are drawn from $P_{\text{target}}$ and $P_{\text{target}}(x,y)\neq P_{i}(x,y)$ for $i\in\{1,2,\dots,M\}$.
To protect privacy, the server is not permitted to directly access the data on the clients. After local training, each client uploads its model parameters to the server, where the parameters are aggregated and then sent back to the clients. Let $\theta_{i}^{t}$ denote the model parameters of client $i$ at round $t$. The server aggregates the parameters as follows:
$\theta^{t+1}=\frac{1}{N}\sum_{i=1}^{K\cdot M}n_{i}\theta_{i}^{t},$ (2)
where $N=\sum_{i=1}^{K\cdot M}n_{i}$ is the total number of samples across all clients [19] and $n_{i}$ denotes the number of samples on client $i$. After aggregation, the server sends the updated parameters $\theta^{t+1}$ back to each client.
3.2 Algorithm Overview
In FedCCRL, in addition to transmitting model parameters between the clients and the server, channel-wise sample statistics are also transmitted.
On the client side, for a batch of samples $X\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ is the number of channels, and $H$ and $W$ represent the height and width of the images respectively, we first utilize feature augmentation to generate two batches of augmented samples:
$X^{(1)}=\mathcal{M}(X),X^{(2)}=\mathcal{M}(X),$ (3)
where $\mathcal{M}(\cdot)=\text{AugMix}\circ\text{MixStyle}$ represents the composition of MixStyle and AugMix. The detailed augmentation processes of MixStyle and AugMix are provided in Algorithm 2 and Algorithm 3, respectively.
After applying data augmentation, we extract representations and generate predictions for each batch, which are then used to compute the overall loss. Let $\mathcal{Z}$ denote the representation space. We formalize the model $f$ as $f=g\circ h$, where $h:\mathcal{X}\rightarrow\mathcal{Z}$ represents the representation encoder and $g:\mathcal{Z}\rightarrow\hat{\mathcal{Y}}$ denotes the classifier. We extract representations and produce predictions as follows:
$Z=h(X),Z^{(1)}=h(X^{(1)}),Z^{(2)}=h(X^{(2)}),$ (4)
$\hat{Y}=g(Z),\hat{Y}^{(1)}=g(Z^{(1)}),\hat{Y}^{(2)}=g(Z^{(2)}).$ (5)
We then employ supervised contrastive loss for representation alignment, as outlined in Equation 12, and apply Jensen-Shannon divergence for prediction alignment as specified in Equation 14, to ensure consistent predictions between the original and generated samples.
Algorithm 1 FedCCRL
Input: Ratio of statistics to be uploaded $r\in(0,1)$, Hyper-parameters $\lambda_{1}$ and $\lambda_{2}$, Communication rounds $T$, Training epochs $E$.
Output: Model parameter $\theta$
1:Server initializes $f$ and distributes it to each client.
2:for each round $t=1,2,\ldots,T$ do
3:     for each client $i=1,2,\ldots,KM$ do
4:         $F_{i}\leftarrow$ Eq. 8
5:         Client $i$ uploads $F_{i}$ to the server
6:     end for
7:     Server constructs $F_{pool}$ and distributes $F_{pool}^{i}$ to each client $i$
8:     for each client $i=1,2,\ldots,KM$ do
9:         for each epoch $e=1,2,\ldots,E$ do
10:              for each batch $X$ in $D_{i}$ do
11:                  $X^{(1)},X^{(2)}$$\leftarrow$ Eq. 3 $\triangleright$ Algorithm 2, 3
12:                  $Z,Z^{(1)},Z^{(2)}$$\leftarrow$ Eq. 4
13:                  $\hat{Y},\hat{Y}^{(1)},\hat{Y}^{(2)}$$\leftarrow$ Eq. 5
14:                  $\mathcal{L}_{RA}\leftarrow$ Eq. 12
15:                  $\mathcal{L}_{JS}\leftarrow$ Eq. 14
16:                  $\mathcal{L}_{CLS}\leftarrow$ Eq. 16
17:                  $\mathcal{L}\leftarrow$ Eq. 15
18:                  $\theta_{i}^{t}\leftarrow$ update model with Adam
19:              end for
20:         end for
21:         Client $i$ sends $\theta_{i}^{t}$ to the server
22:     end for
23:     $\theta^{t+1}\leftarrow$ Eq. 2
24:end for
25:return $\theta^{T+1}$
3.3 Feature Augmentation
In the context of DG, we identify two types of features:
  • Domain-specific features: These are features that change across domains such as color distribution, lighting conditions, or style, which have little to do with the label of the image.
  • Domain-invariant features: These features, including contours, textures, and shapes, are directly linked to the image’s label and remain consistent across different domains.
The core idea behind the feature augmentation component in FedCCRL is to perform cross-client style transfer on domain-specific features and to perturb domain-invariant features. To achieve this, we adapt MixStyle to the FDG setting and combine it with AugMix. Specifically, MixStyle handles the cross-client style transfer, while AugMix is responsible for perturbing the domain-invariant features.
MixStyle operates by leveraging the channel-wise statistics of samples. Given a sample $x\in\mathbb{R}^{C\times H\times W}$, we compute the channel-wise mean and standard deviation as follows:
$\mu(x)_{c}=\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}x_{c,h,w}\text{,}$ (6)
$\sigma(x)_{c}=\sqrt{\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}(x_{c,h,w}-\mu(x)_{c})^{2}}\text{.}$ (7)
In FedCCRL, in addition to transporting model parameters to the server, clients are also required to upload sample statistics. Since attackers cannot reconstruct the original samples using only channel-wise means and standard deviations, our method effectively safeguards users’ data privacy.
Specifically, each client $i$ uploads a certain proportion of sample statistics denoted as:
$F_{i}=\{(\mu(x_{j}),\sigma(x_{j}))\mid x_{j}\in D_{i},j=1,2,\ldots,\lceil rn_{i}\rceil\},$ (8)
where $r\in(0,1)$ represents the ratio of statistics to be uploaded.
After collecting statistics from all clients, the server concatenates these statistics to get the pool of statistics: $F_{\text{pool}}=\{F_{i}\mid i=1,2,\ldots,K\cdot M\}$, and subsequently distributes $F_{pool}^{i}=F_{pool}\setminus F_{i}$ to each client $i$. Each client apply MixStyle to generate style transferred samples according to Algorithm 2.
Algorithm 2 MixStyle on Client $i$
Input: A batch of samples $X$, Pool of statistics $F_{pool}^{i}$, $\alpha\in(0,\infty)$
Output: A batch of style transferred samples $X_{mixstyle}$
1:$\hat{X}=[\ ]$
2:for $x$ in $X$ do
3:     Compute $\mu(x)$, $\sigma(x)$triangleright Equation 6, 7
4:     Normalize the input sample: $\tilde{x}=\frac{x-\mu(x)}{\sigma(x)}$
5:     Uniformly sample a pair of statistics $(\mu^{\prime},\sigma^{\prime})\sim F_{pool}$
6:     Sample $\lambda\sim\text{Beta}(\alpha,\alpha)$
7:     $\gamma_{mix}=\lambda\mu^{\prime}+(1-\lambda)\mu(x)$
8:     $\beta_{mix}=\lambda\sigma^{\prime}+(1-\lambda)\sigma(x)$
9:     $\hat{x}=\gamma_{mix}\tilde{x}+\beta_{mix}$
10:     $\hat{X}$.append($\hat{x}$)
11:end for
12:$X_{mixstyle}=\text{cat}(\hat{X})$
Notably, the computational and transmission overhead associated with sample statistics is negligible compared to the costs of model training and transmission. Besides, as shown in Figure 6, our method achieves better performance than other baselines even with small upload ratio $r$. Therefore, FedCCRL does not introduce significant computational or communication overhead.
In addition to MixStyle, we employ AugMix [8] to perturb domain-invariant features. AugMix takes the original sample $x$ and processes it through multiple augmentation chains, each consisting of a sequence of simple augmentation operations. The outputs from these chains are then mixed using an element-wise convex combination. The detailed process of AugMix is provided in Algorithm 3 in supplementary material, and we adopt the same set of operations $\mathcal{O}$ as described in [8].
In Figure 3, we present the feature augmentation effects of MixStyle and AugMix. It is evident that MixStyle transfers domain-specific features, such as color distribution, while AugMix effectively perturbs domain-invariant features, such as contours of objects.
(a) Original Image
(b) MisStyle
(c) AugMix
Figure 3: Effect of feature augmentation. (b) is generated from (a) using MixStyle, and (c) is generated from (b) by AugMix.
3.4 Representation Alignment
To enable the model to focus on extracting domain-invariant features that capture the essential information relevant to the task, we employ supervised contrastive loss [12] to align feature representations across original samples and augmented samples. Supervised contrastive loss leverages label information to ensure that representations from the same class are pulled closer together in the latent space, while representations from different classes are pushed apart. This alignment mechanism enables the model to extract correct semantic features.
Given any two batches of representations $Z^{\prime},Z^{\prime\prime}\in\mathbb{R}^{B\times V}$ and their corresponding labels $Y^{\prime}$ and $Y^{\prime\prime}$, we concatenate these two batches of features and labels:
$Z=\text{cat}(Z^{\prime},Z^{\prime\prime})\in\mathbb{R}^{2B\times V},$ (9)
$Y=\text{cat}(Y^{\prime},Y^{\prime\prime}),$ (10)
where $V$ denotes the size of feature representations.
Let $I=\{1,2,\ldots,2B\}$ to be the index set and $A(i)=I\setminus i$. The supervised contrastive loss of $Z^{\prime}$ and $Z^{\prime\prime}$ is calculated as follow:
$\mathcal{L}_{\text{SC}}(Z^{\prime},Z^{\prime\prime})=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{s(Z_{i},Z_{p})}{\sum_{a\in A(i)}s(Z_{i},Z_{a})},$ (11)
where $P(i)=\{p\in A(i)\mid y_{p}=y_{i}\}$, $s(Z_{i},Z_{p})=\exp(\text{sim}(Z_{i},Z_{p})/\tau)$. $\text{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is a temperature parameter that controls the scale of similarities.
The representation alignment loss $\mathcal{L}_{\text{RA}}$ in FedCCRL is defined as:
$\mathcal{L}_{\text{RA}}=\frac{1}{2}(\mathcal{L}_{SC}(Z^{(1)},Z)+\mathcal{L}_{SC}(Z^{(2)},Z)).$ (12)
3.5 Jensen-Shannon Divergence
As $X^{(1)}$ and $X^{(2)}$ are generated from $X$, we expect the model to not only correctly classify the original and augmented samples, but also produce similar predictions $\hat{Y}$, $\hat{Y}^{(1)}$ and $\hat{Y}^{(2)}$. We utilize Jensen-Shannon divergence [20] minimization to achieve consistent predictions across original and generated samples. The Jensen-Shannon divergence is designed to measure the similarity between these probability distributions by averaging them and then computing their divergence from the mean. Specifically, let the average prediction distribution be defined as:
$\bar{Y}=\frac{1}{3}(\hat{Y}+\hat{Y}^{(1)}+\hat{Y}^{(2)}).$ (13)
The Jensen-Shannon divergence is then computed as follows:
$\mathcal{L}_{JS}=\frac{1}{3}(\text{KL}(\hat{Y},\bar{Y})+\text{KL}(\hat{Y}^{(1)},\bar{Y})+\text{KL}(\hat{Y}^{(2)},\bar{Y}))$ (14)
where $\text{KL}(\cdot,\cdot)$ denotes KL divergence.
3.6 Overall Loss Function
The overall loss function is formulated as follows:
$\mathcal{L}=\mathcal{L}_{CLS}+\lambda_{1}\mathcal{L}_{RA}+\lambda_{2}\mathcal{L}_{JS},$ (15)
where the classification loss is computed as:
$\mathcal{L}_{CLS}=\sum_{Y^{\prime}\in\{\hat{Y},\hat{Y}^{(1)},\hat{Y}^{(2)}\}}\frac{1}{3}\mathcal{L}_{CRE}(Y^{\prime},Y),$ (16)
and $\mathcal{L}_{CRE}$ denotes the cross-entropy loss. The terms $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the relative importance of the representation alignment loss $\mathcal{L}_{RA}$ and the Jensen-Shannon divergence loss $\mathcal{L}_{JS}$, respectively.
4 Experiments
PACS OfficeHome miniDomainNet
P A C S Avg. A C P R Avg. C P R S Avg.
FedAvg 90.48 69.19 76.15 71.62 76.86 62.30 49.51 76.12 77.55 66.37 63.90 57.64 67.90 49.63 59.77
FedProx 90.60 69.09 74.32 70.15 76.04 63.08 50.13 76.28 77.65 66.78 63.74 57.65 68.02 49.81 59.81
FedADG 91.02 65.04 72.95 65.82 73.71 60.82 47.93 71.93 75.58 64.06 54.79 52.32 62.37 41.60 52.77
GA 92.81 66.50 76.19 69.56 76.26 61.80 49.60 75.51 77.09 66.00 63.80 57.37 67.62 49.83 59.66
FedSR 91.74 72.36 75.55 70.48 77.53 64.19 50.29 76.39 77.65 67.13 62.98 56.79 67.35 49.39 59.13
FedIIR 91.26 71.73 77.94 71.32 78.06 61.48 49.97 75.40 77.25 66.03 63.90 57.11 67.90 49.33 59.56
CCST 90.16 76.37 76.11 78.58 80.30 61.20 50.48 76.53 77.48 66.42 62.75 56.54 65.32 47.93 58.14
FedCCRL 93.83 79.15 77.9 78.95 82.46 63.08 55.67 76.77 77.71 68.31 66.33 59.68 68.03 54.50 62.14
Table 1: Test accuracy on each dataset. In this set of experiments, we set the upload ratio of the statistics $r=0.1$. To ensure fairness, each algorithm is evaluated three times, and the final average is taken as the experimental result.
Figure 4: Average test accuracy (%) versus the number of participating clients.
4.1 Dataset and Evaluation Protocol
Our algorithm and baselines are evaluated on three widely used datasets in DG: PACS [14], OfficeHome [30], and miniDomainNet [42]. PACS comprises 9,991 samples across four domains: Art Painting, Cartoon, Photo, and Sketch, and is categorized into seven classes. OfficeHome contains 15,588 samples from four domains: Art, Clipart, Product, and Real World, and is divided into 65 classes. miniDomainNet features 140,006 samples from four domains: Clipart, Infograph, Painting, and Real, and encompasses a total of 126 classes.
We adopt the leave-one-domain-out evaluation protocol [14], where one domain is held out as the test set while the remaining domains are used for training. This process is repeated for each domain to ensure robust evaluation across varying distribution shifts.
4.2 Baselines and Implementation Details
In this experiment, we compare FedCCRL with several traditional FL frameworks, including FedAvg [19] and FedProx [16], as well as state-of-the-art FDG approaches such as FedADG [39], FedSR [24], FedIIR [7], GA [40], and CCST [3].
Following the data partition scheme outlined in Section 3.1, we evaluate FedCCRL and the baselines across different numbers of clients. We utilize MobileNetV3-Large [9] as the backbone network, where the final fully connected layer acts as the classifier $g$, and the preceding layers serve as the representation encoder $h$.
For training, we set the number of communication rounds to 10, and the number of local training epochs is fixed at 3 for all methods. The Adam optimizer [13] is employed with an initial learning rate of 0.001. Furthermore, MixStyle is configured with the parameter $\alpha=0.1$, while for AugMix, the parameter $\beta$ is set to 1.0. For supervised contrastive loss, we set the temperature parameter $\tau$ to 0.1. Input samples from the PACS and OfficeHome datasets are resized to $224\times 224$, whereas those from miniDomainNet are resized to $128\times 128$. To manage the learning rate, we adopt a cosine learning rate scheduler, which facilitates a smooth decay of the learning rate throughout the training process.
To guarantee the reliability of our proposed algorithms, we run each experiment three times independently and report the averaged results.
4.3 Comparison with Baselines
We evaluate FedCCRL and other baselines under scenarios with different numbers of clients. Table 1 specifically shows the performance of each algorithm when the total number of clients is set to 6. In this table, we use the first letter of each domain to represent the domain, and each letter column corresponds to the performance when that domain is used as the target domain. The ”Avg.” column provides the average results across the four leave-one-domain-out experiments. Additionally, Figure 4 presents the performance of each algorithm under varying client numbers, where the results are averaged across the four leave-one-domain-out experiments. In these experiments, the upload ratio for FedCCRL is set to 0.1.
From Table 1, we have the following observations:
  • FedCCRL outperforms all baseline methods on all datasets. Specifically, FedCCRL achieves the best average accuracy on all datasets. In addition, FedCCRL attains top scores on each test domain of both PACS and miniDomainNet datasets.
  • Most baselines specifically designed for FDG fail to achieve consistent performance across different datasets. In some cases, their performance is even inferior to traditional federated learning algorithms such as FedAvg and FedProx.
Additionally, as shown in Figure 4, although the performance of all baselines significantly declines as the number of participating clients increases, FedCCRL consistently outperforms the baselines across both small- and large-scale client scenarios.
4.4 Ablation Study
MixStyle AugMix $\mathcal{L}_{RA}$ $\mathcal{L}_{JS}$ PACS OfficeHome miniDomainNet
79.92 66.58 59.57
80.95 66.93 61.43
80.52 67.34 60.82
81.11 67.58 61.83
79.32 66.84 59.64
80.62 67.29 60.94
80.38 67.68 61.53
80.97 67.92 61.94
80.48 67.29 60.40
81.18 67.81 61.47
80.90 68.08 61.49
82.46 68.31 62.14
Table 2: Ablation study results. The total number of clients is set to 6 and the upload ratio is set to 0.1.
We investigate the effects of MixStyle, AugMix, representation alignment regularizer $\mathcal{L}_{RA}$, and Jensen-Shannon divergence $\mathcal{L}_{JS}$ on FedCCRL across the PACS, OfficeHome, and miniDomainNet datasets. The results are presented in Table 2, from which we can draw the following conclusions:
  • All four components significantly enhance model generalization. Using representation alignment loss or Jensen-Shannon divergence individually outperforms cross-entropy alone, and their combination yields the best performance. Additionally, combining AugMix and MixStyle improves results compared to using either technique alone.
  • Even without regularization, FedCCRL with only data augmentation techniques outperforms all baselines. As shown in Table 1, the combination of MixStyle and AugMix surpasses all baselines across the three datasets.
Figure 5: t-SNE visualization of representations of each test domain on PACS dataset.
4.5 Impact of Upload Ratio and Hyperparameters
Figure 6: Effect of $r$ and $\lambda_{1}$, $\lambda_{2}$ on accuracy across datasets
Figure 7: Grad-CAM visualization on the PACS dataset.
We evaluated the performance of FedCCRL under different upload ratios and hyperparameter settings, and the results are presented in Figure 6. When evaluating the effect of $\lambda_{1}$, $\lambda_{2}$ is fixed at 1. Similarly, when assessing the impact of $\lambda_{2}$, $\lambda_{1}$ is set to 0.1.
The evaluation results in Figure 6 demonstrate that both the upload ratio $r$ and the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ have minimal impact on the overall performance across the PACS, OfficeHome, and miniDomainNet datasets. Therefore, it can be concluded that FedCCRL is robust to variations in these parameters and outperforms other baselines even with only a very small amount of upload of statistics.
4.6 Feature Distribution
To better illustrate the effect of FedCCRL on the model, Figure 5 and Figure 7 respectively provide the t-SNE visualization of representation distribution and Grad-CAM visualization of the model’s concentration on the PACS dataset.
Figure 5 demonstrates that FedCCRL achieves better separation of classes compared to FedAvg, indicating that FedCCRL is effective in learning clearer, domain-invariant feature representations across different domains, reducing overlap between the representations of different classes. Additionally, Figure 7 reveals that the model trained with FedCCRL focuses more effectively on domain-invariant features, such as the giraffe’s head and the elephant’s trunk. This suggests that FedCCRL enhances the model’s ability to concentrate on relevant, discriminative regions in the images, leading to improved generalization across domains.
5 Conclusion
In this paper we propose FedCCRL, a federated domain generalization framework that improves model generalization across unseen domains without violating privacy rules or introducing excessive computational and communication costs. Specifically, FedCCRL effectively combines MixStyle and AugMix for cross-client style transfer and the perturbation of domain-invariant features. In addition, FedCCRL utilizes supervised contrastive loss for representation alignment and apply Jensen-Shannon divergence for prediction alignment. Experiments on PACS, OfficeHome, and miniDomainNet demonstrate that FedCCRL outperforms existing baselines, demonstrating its effectiveness in enhancing models’ generalization capability.
Supplementary Material
6 Process of AugMix
Algorithm 3 AugMix
Input: A batch of samples $X$, Operations base $\mathcal{O}$, $\beta\in(0,\infty)$
Output: A batch of augmented samples $X_{augmix}$
1:$\hat{X}=[\ ]$
2:for $x$ in $X$ do
3:       Fill $x_{\text{aug}}$ with zeros
4:       Uniformly sample $k\sim\{1,2,3\}$
5:       Sample $(w_{1},w_{2},\ldots,w_{k})\sim\text{Dirichlet}(\beta,\beta,\ldots,\beta)$
6:       for $i=1,\ldots,k$ do
7:             Sample operations $\text{op}_{1},\text{op}_{2},\text{op}_{3}\sim\mathcal{O}$
8:             $\text{op}_{12}=\text{op}_{2}\circ\text{op}_{1}$, $\text{op}_{123}=\text{op}_{3}\circ\text{op}_{2}\circ\text{op}_{1}$
9:             Uniformly sample $\text{chain}\sim\{\text{op}_{1},\text{op}_{12},\text{op}_{123}\}$
10:             $x_{\text{aug}}\mathrel{+}=w_{i}\cdot\text{chain}(x)$triangleright Element-wise Addition
11:       end for
12:       Sample weight $m\sim\text{Beta}(\beta,\beta)$
13:       $\hat{x}=mx+(1-m)x_{\text{aug}}$
14:       $\hat{X}.\text{append}(\hat{x})$
15:end for
16:$X_{augmix}=\text{cat}(\hat{X})$